# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 2.3 Problem setup
Let $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a strongly convex objective function. Recall the following non-stationary stochastic optimization problem:
$\bm{\theta}_{t}^{\star}\in\operatorname*{arg\,min}_{\bm{\theta}\in\mathbb{R}^{d}}\mathbb{E}_{X_{t}\sim\Pi_{t}}[g(\bm{\theta},X_{t})].$
Here $g(\bm{\theta},X_{t})$ is a noise perturbed measurement of $G(\bm{\theta})$ and $X_{t}$ is a random variable sampled from a distribution $\Pi_{t}$. Let $(\mathcal{F}_{t})_{t\geq 1}$ be the natural filtration $\mathcal{F}_{t}=\sigma(X_{0},\dots,X_{t})$. Throughout the entirety of our analysis, we will be imposing an assumption modeling stochasticity in the non-stationary setting:
@@
There exists a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ with $\mathcal{F}_{0}=\{\emptyset,\Omega\}$. Let $(X_{t})_{t\geq 0}$ be an $\mathbb{F}$–adapted process, i.e., $X_{t}$ is $\mathcal{F}_{t}$–measurable for all $t$. For each $t\geq 0$, define the conditional risk $F_{t+1}(\bm{\theta}):=\mathbb{E}\big[g(\bm{\theta},X_{t+1})\mid\mathcal{F}_{t}\big]$, and let $\bm{\theta}_{t+1}^{\star}$ denote its (a.s. unique) minimizer: $\bm{\theta}_{t+1}^{\star}\in\operatorname*{arg\,min}_{\bm{\theta}\in\mathbb{R}^{d}}F_{t+1}(\bm{\theta})$. Assume the following hold for all $t\geq 0$:
  1. i)
    (Predictable target) $\bm{\theta}_{t+1}^{\star}$ is $\mathcal{F}_{t}$–measurable.
  2. ii)
@@
$\begin{split}\bm{\theta}_{t+1}&=\bm{\theta}_{t}-\gamma_{t}\nabla_{\theta}g(\bm{\theta}_{t},X_{t+1})\\ &=\bm{\theta}_{t}-\gamma_{t}m_{t+1}(\bm{\theta}_{t})-\gamma_{t}\xi_{t+1}(\bm{\theta}_{t}).\end{split}$ (SGD)
@@
$\begin{split}&\bm{\psi}_{t}=\bm{\theta}_{t}+\beta_{1}(\bm{\theta}_{t}-\bm{\theta}_{t-1})\\ &\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}\nabla_{\bm{\theta}}g(\bm{\psi}_{t},X_{t+1})+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).\end{split}$ (2.3)
When $\beta_{1}=0$ and $\beta_{2}=\beta$, then we recover Polyak’s heavy-ball method. If $\beta_{2}=0$ and $\beta_{1}=\beta$, we recover Nesterov’s accelerated method. Thus to capture both these methods simultaneously, we assume $\beta_{1}+\beta_{2}=\beta$ and $\beta_{1}\beta_{2}=0$ for some constant fixed $\beta\in[0,1)$. We adapt this formulation from [YYS16]. Using this, our parameter updates can be written as follows:
$\begin{split}&\bm{\psi}_{t}=\bm{\theta}_{t}+\beta_{1}(\bm{\theta}_{t}-\bm{\theta}_{t-1})\\ &\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}m_{t+1}(\bm{\psi}_{t})-\gamma_{t}\xi_{t+1}(\bm{\psi}_{t})+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).\end{split}$ (SGDM)
Note in this setup that $\bm{\theta}_{t+1}$ is not $\mathcal{F}_{t}$ measurable. Rather if we define $\mathcal{F}\subset\mathcal{F}_{t}^{+}=\sigma(X_{0},\dots,X_{t},X_{t+1})$, then $\bm{\theta}_{t+1}$ is $\mathcal{F}_{t}^{+}$ measurable. In order to examine the convergence of standard and momentum based stochastic gradient methods, we will need to make some assumptions on the conditional mean gradient map $\bm{\theta}\mapsto m_{t}(\bm{\theta})$. This amounts to assuming that the map $\bm{\theta}\mapsto m_{t}(\bm{\theta})$ is uniformly $\mu$-strongly monotone and Lipschitz. Note that this is equivalent to assuming that $\mathbb{E}_{X_{t}\sim\Pi_{t}}[g(\bm{\theta},X_{t})]$ is $\mu$-strongly convex and has Lipschitz gradients (under mild technical conditions for which the Dominated Convergence Theorem holds). Under strong convexity, the minimizer $\bm{\theta}_{t}^{\star}$ is unique. It should be noted that these conditions are standard in the analysis of SGD and SGD with momentum in the literature [YYS16] and are satisfied by many problems of interest, especially when regularization is used (e.g. mean-square error risks, logistic risks, etc).
@@ 3.1 Bounds in expectation
In this section, we will obtain guarantees on the tracking error $\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|$ in expectation. The intermediate steps used to get this bound will be useful to obtain high-probability bounds that we discuss later. Before proceeding, we will make a few assumptions regarding the second moments of the minimizer drift $\Delta_{t}=\|\bm{\theta}_{t}^{\star}-\bm{\theta}_{t+1}^{\star}\|$ and the gradient noise, simply for presentation. These assumptions are not essential for the result, but we adopt them to simplify the exposition.
Assumption 3.1 (Conditional second-moment bounds).
@@
  1. i)
    (Minimizer drift) The minimizer drift $\Delta_{t}$ satisfies $\mathbb{E}\big[\Delta_{t}^{2}\mid\mathcal{F}_{t}\big]\leq\Delta^{2}\;\;\text{a.s.}$
  2. ii)
@@
    (Gradient noise along iterates) The gradient noise $\bm{\xi}_{t+1}$ satisfies $\mathbb{E}\big[\|\bm{\xi}_{t+1}(\bm{\theta}_{t})\|^{2}\mid\mathcal{F}_{t}\big]\leq\sigma^{2}\;\;\text{a.s.}$
@@
We can now establish the following theorem that establishes the expected tracking error for SGD in nonstationary time-series environments. We defer the proofs for this section to Section˜A.1.
Theorem 3.1 (Tracking error bound in expectation for SGD).
@@
Under ˜3.1, for all $t\geq 0$ and $\gamma\leq\min\left\{\mu/L^{2},1/L\right\}$, the following tracking error bound holds in expectation for SGD:
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\left(1-\frac{\gamma\mu}{2}\right)^{t}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{4\Delta^{2}}{\gamma^{2}\mu^{2}}+\frac{\sigma^{2}\gamma}{\mu}.$
@@
This bound is similar to [CDH21] as it consists of a contraction term that arises from optimization which decays linearly in $t$, a drift/tracking term that depends on the minimizer drift, and a noise term. These contributions are irreducible for constant stepsize: even as $t\to\infty$, SGD cannot converge arbitrarily close to $\bm{\theta}_{t}^{\star}$ because (i) the optimizer itself moves over time (drift), and (ii) stochastic gradients inject persistent variance (noise). Moreover, the two steady-state terms exhibit an explicit stepsize tradeoff: larger $\gamma$ amplifies the noise floor, while smaller $\gamma$ reduces the noise but worsens tracking of a moving target. Consequently, we can show that after a burn-in period which varies based on whether we use a constant stepsize or epoch-wise step-decay schedule, we can reach the irreducible floor of order $\mathcal{O}\left(\sigma^{2}\gamma/\mu+\Delta^{2}/\gamma^{2}\mu^{2}\right)$.

### v3

@@ 2.1 Problem setup
Recall the optimization problem (Opt). Let $(\mathcal{F}_{t})_{t\geq 1}$ be the natural filtration $\mathcal{F}_{t}=\sigma(X_{0},\dots,X_{t})$. Throughout the entirety of our analysis, we will be imposing an assumption modeling stochasticity in the non-stationary setting:
@@
There exists a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ with $\mathcal{F}_{0}=\{\emptyset,\Omega\}$. Let $(X_{t})_{t\geq 0}$ be an $\mathbb{F}$–adapted process, i.e., $X_{t}$ is $\mathcal{F}_{t}$–measurable for all $t$. For each $t\geq 0$, let $\Pi_{t+1}$ denote the regular conditional law of $X_{t+1}$ given $\mathcal{F}_{t}$, i.e., $\Pi_{t+1}(A)=\mathbb{P}(X_{t+1}\in A\mid\mathcal{F}_{t})$ a.s. for every measurable set $A$, and assume $\Pi_{t+1}$ is $\mathcal{F}_{t}$–measurable. Define the conditional risk
$G_{t+1}(\bm{\theta}):=\mathbb{E}\!\left[g(\bm{\theta},X_{t+1})\mid\mathcal{F}_{t}\right]=\mathbb{E}_{X\sim\Pi_{t+1}}\!\left[g(\bm{\theta},X)\right],$
and let $\bm{\theta}_{t+1}^{\star}\in\arg\min_{\bm{\theta}\in\mathbb{R}^{d}}G_{t+1}(\bm{\theta})$ denote a (measurable) minimizer. Assume the following hold for all $t\geq 0$:
  1. 1.
    (Predictable minimizer) $\bm{\theta}_{t+1}^{\star}$ is $\mathcal{F}_{t}$–measurable. [^1] Since $G_{t+1}(\bm{\theta})=\mathbb{E}[g(\bm{\theta},X_{t+1})\mid\mathcal{F}_{t}]$ is $\mathcal{F}_{t}$–measurable for each fixed $\bm{\theta}$, standard measurable selection conditions ensure existence of an $\mathcal{F}_{t}$–measurable minimizer (e.g., Carathéodory regularity plus compactness/coercivity yielding a.s. nonempty $\arg\min$). If the minimizer is a.s. unique (e.g., under strong convexity), measurability follows by the measurable maximum theorem.
  2. 2.
@@
$\begin{split}\bm{\theta}_{t+1}&=\bm{\theta}_{t}-\gamma_{t}\nabla_{\theta}g(\bm{\theta}_{t},X_{t+1})\\ &=\bm{\theta}_{t}-\gamma_{t}m_{t+1}(\bm{\theta}_{t})-\gamma_{t}\bm{\xi}_{t+1}(\bm{\theta}_{t}).\end{split}$ (SGD)
@@
$\begin{split}&\bm{\psi}_{t}=\bm{\theta}_{t}+\beta_{1}(\bm{\theta}_{t}-\bm{\theta}_{t-1})\\ &\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}\nabla_{\bm{\theta}}g(\bm{\psi}_{t},X_{t+1})+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).\end{split}$
When $\beta_{1}=0$ and $\beta_{2}=\beta$, we recover Polyak’s Heavy-Ball method. If $\beta_{2}=0$ and $\beta_{1}=\beta$, we recover Nesterov’s accelerated method. Thus to capture both these methods simultaneously, we assume $\beta_{1}+\beta_{2}=\beta$ and $\beta_{1}\beta_{2}=0$ for some constant fixed $\beta\in[0,1)$. We adopt this formulation from (Yuan et al., 2016). Using this, our parameter updates can be written as follows:
$\begin{split}&\bm{\psi}_{t}=\bm{\theta}_{t}+\beta_{1}(\bm{\theta}_{t}-\bm{\theta}_{t-1})\\ &\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}m_{t+1}(\bm{\psi}_{t})-\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ &+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).\end{split}$ (SGDM)
Note that $\bm{\theta}_{t+1}$ is $\mathcal{F}_{t+1}$-measurable rather than $\mathcal{F}_{t}$-measurable. To analyze the convergence of (SGD) and (SGDM), we assume the conditional mean gradient map $\bm{\theta}\mapsto m_{t}(\bm{\theta})$ is uniformly $\mu$-strongly monotone and $L$-Lipschitz. Under mild regularity conditions (e.g. sufficient integrability for the Dominated Convergence Theorem), this is equivalent to assuming $G_{t}(\bm{\theta})$ is $\mu$-strongly convex and $L$-smooth, which ensures that the minimizer $\bm{\theta}_{t}^{\star}$ is unique.
@@
These assumptions are standard in the SGD literature (Yuan et al., 2016) and are satisfied by many problems of interest, especially when regularization is used (e.g. mean-square error risks, logistic risks, etc). We also define the condition number as $\kappa:=L/\mu$.
@@ 3.1 Bounds in expectation
In this section, we will obtain guarantees on the tracking error $\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|$ in expectation. The intermediate steps used to obtain this bound will be useful in obtaining high-probability bounds, which we discuss later. Before proceeding, we will make a few assumptions regarding the second moments of the minimizer drift $\Delta_{t}=\|\bm{\theta}_{t}^{\star}-\bm{\theta}_{t+1}^{\star}\|$ and the gradient noise, simply for presentation. These assumptions are not essential for the result, but we adopt them to simplify the exposition.
Assumption 3.1 (Second-moment bounds).
@@
  1. 1.
    (Minimizer drift) The minimizer drift $\Delta_{t}$ satisfies $\mathbb{E}\big[\Delta_{t}^{2}\big]\leq\Delta^{2}\;\;\text{a.s.}$
  2. 2.
@@
    (Gradient noise along iterates) The gradient noise $\bm{\xi}_{t+1}$ satisfies $\mathbb{E}\big[\|\bm{\xi}_{t+1}(\bm{\theta}_{t})\|^{2}\big]\leq\sigma^{2}$ and $\mathbb{E}\big[\|\bm{\xi}_{t+1}(\bm{\psi}_{t})\|^{2}\big]\leq\sigma^{2}\;\;\text{a.s.}$
@@
We can now state the following theorem that establishes the expected tracking error for (SGD) in nonstationary stochastic environments. We defer the proofs for this section to Appendix C.
Theorem 3.1 (Tracking error bound in expectation for (SGD)).
@@
Under Assumption 3.1, $\forall t\geq 0$ and $\gamma\leq\min\left\{\mu/L^{2},1/L\right\}$, the following tracking error bound holds in expectation for (SGD):
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\left(1-\frac{\gamma\mu}{2}\right)^{t+1}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{4\Delta^{2}}{\gamma^{2}\mu^{2}}+\frac{\sigma^{2}\gamma}{\mu}.$
@@
This bound is similar to Cutler et al. (2021), as it consists of a contraction term that arises from optimization that decays linearly in $t$, a drift/tracking term that depends on the minimizer drift, and a noise term. These contributions are irreducible for constant stepsize: even as $t\to\infty$, (SGD) cannot converge arbitrarily close to $\bm{\theta}_{t}^{\star}$ because (i) the optimizer itself moves over time (drift), and (ii) stochastic gradients inject persistent variance (noise). Moreover, the two steady-state terms exhibit an explicit stepsize tradeoff: larger $\gamma$ amplifies the noise floor, while smaller $\gamma$ reduces the noise but worsens tracking of a moving minima. Consequently, we can show that after a burn-in period that varies depending on whether we use a constant stepsize or an epoch-wise step-decay schedule, we can reach the irreducible floor of order $\mathcal{O}\left(\sigma^{2}\gamma/\mu+\Delta^{2}/\gamma^{2}\mu^{2}\right)$.
Theorem 3.2 (Time to reach the asymptotic tracking error in expectation for (SGD)).
$\mathcal{E}(\gamma):=\frac{\sigma^{2}\gamma}{\mu}+\frac{4\Delta^{2}}{\mu^{2}\gamma^{2}},\;\gamma^{\star}\in\arg\min_{\gamma\in(0,1/2L]}\mathcal{E}(\gamma)$
with $\mathcal{E}:=\mathcal{E}(\gamma^{\star})$. Then we have the following:
  1. 1.
    (Constant learning rate). If $\gamma_{t}\equiv\gamma^{\star}$, then for all $t\geq 0$, $\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\lesssim\mathcal{E}$ after time $t\lesssim(1/\mu\gamma^{\star})\log\Bigl(\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}/\mathcal{E}\Bigr).$
  2. 2.
    $\gamma^{\star}=\Bigl(\frac{8\Delta^{2}}{\mu\sigma^{2}}\Bigr)^{1/3},\;\mathcal{E}=3\Bigl(\frac{\Delta\sigma^{2}}{\mu^{2}}\Bigr)^{2/3}.$
    Define epochs $k=0,1,\dots,K-1$ with $\gamma_{0}:=1/2L$
    $\gamma_{k}:=\frac{\gamma_{k-1}+\gamma^{\star}}{2},\;K:=1+\Bigl\lceil\log_{2}\Bigl(\frac{\gamma_{0}}{\gamma^{\star}}\Bigr)\Bigr\rceil,$
    and epoch lengths
    $T_{0}:=\Bigl\lceil\frac{2}{\mu\gamma_{0}}\log\Bigl(\frac{2\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}}{\mathcal{E}(\gamma_{0})}\Bigr)\Bigr\rceil,\;T_{k}:=\Bigl\lceil\frac{2\log 4}{\mu\gamma_{k}}\Bigr\rceil.$
    Run (SGD) with constant stepsize $\gamma_{k}$ for $T_{k}$ iterations in epoch $k$, starting from $\theta_{0}$. Let $T:=\sum_{k=0}^{K-1}T_{k}$ be the total horizon. Then the final iterate satisfies $\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\lesssim\mathcal{E}\;\text{after time}\;t\lesssim(L/\mu)\cdot\log\Bigl(\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}/\mathcal{E}\Bigr)\;+\;\sigma^{2}/\mu^{2}\mathcal{E}$.
Theorem˜3.1 and Theorem˜3.2 provide algorithmic guarantees for (SGD) in non-stationary environments, closely matching those of Cutler et al. (2021) and the static setting (Lan, 2011). Extending these results to (SGDM) is substantially more challenging as the velocity $\bm{v}_{t}$ and position $\bm{\psi}_{t}$ iterates are tightly coupled, precluding reduction to a single recursion that is amenable to standard contraction arguments. Instead, we analyze (SGDM) as a 2D dynamical system on extended state vectors via a Lyapunov function capturing both components. Due to space constraints, we defer the full statement and proof to Appendix C.2.
With this view of (SGDM) as a 2D dynamical system on extended state vectors, we obtain an expectation bound on the tracking error for (SGDM). One will note that we incur extra scaling factors of $(1-\beta)^{-1}$ and $(1-\beta)^{-2}$ which explains why momentum can be significantly worse than (SGD) in non-stationary environments.
Theorem 3.3 (Tracking error bound in expectation for (SGDM)).
Let ˜2.2, ˜2.3, and ˜3.1 hold. Consider the momentum stochastic gradient method (SGDM) and the extended 2D recursion (13). Then when the step-sizes $\gamma$ satisfies $\gamma\leq\mu(1-\beta)^{2}/4L^{2}$, the following tracking error bound holds in expectation for (SGDM):
$\begin{split}&\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{4}{(1-\beta)^{2}}\rho^{2(t+1)}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}\\ &+\frac{48(1+\beta+\gamma L)^{2}}{(1-\beta)^{2}}\cdot\frac{\Delta^{2}}{(1-\rho)^{2}}+\frac{48}{(1-\beta)^{2}}\cdot\frac{\sigma^{2}\gamma^{2}}{1-\rho^{2}}.\end{split}$
In particular taking $\rho=1-\frac{\gamma\mu}{2(1-\beta)}$ and letting $c_{\gamma,\beta}=\exp\!\left(-\gamma\mu/(1-\beta)\right)$, we obtain:
$\begin{split}&\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}c_{\gamma,\beta}^{t+1}\,\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}\\ &+\frac{192(2+\beta)^{2}}{\gamma^{2}\mu^{2}}\Delta^{2}+\frac{96\sigma^{2}\gamma}{\mu(1-\beta)}.\end{split}$
Theorem 3.3 yields the same three-term structure as (SGD): an exponentially decaying contraction term plus irreducible drift and noise-induced floors, but with explicit momentum penalties. In particular, momentum increases sensitivity to initialization via a $(1-\beta)^{-2}$ prefactor, inflates the steady-state noise floor by $(1-\beta)^{-1}$, and slows the effective decay rate to $\exp({-\gamma\mu(t+1)/(1-\beta)})$. Under standard stability tuning, this implies $\gamma\mu\asymp 1/\kappa$ so ill-conditioned problems can exhibit a long burn-in before reaching the steady-state regime. This is especially relevant in deep networks with highly ill-conditioned local curvature (small effective $\mu$) (Sagun et al., 2017; Jastrzębski et al., 2019; Ghorbani et al., 2019; Papyan, 2020). While Nesterov acceleration improves deterministic gradient descent from $\mathcal{O}(\kappa)$ to $\mathcal{O}(\sqrt{\kappa})$ iterations (Nesterov, 1983, 2014), in drifting stochastic settings the dominant error is governed by tracking and variance floors rather than bias decay. Thus the classical $\sqrt{\kappa}$ effect may never materialize, and large $\beta$ primarily raises the noise floor and prolongs burn-in (Theorem˜3.4).